# FENet: Focusing Enhanced Network for Lane Detection

### v3

@@ Abstract
Index Terms—  Lane Detection, Neural Networks, Focusing, Sampling, Self-driving
@@ 1 Introduction
The present study delineates the disparity between human visual focus during driving and the perspectives captured by 2D cameras. Fig. 1 illustrates that experienced drivers preferentially gaze at distant road regions, a strategy crucial for anticipating path geometry and steering adjustments [1, 2, 3]. This behaviour, particularly evident around curves, involves gazing 1-2 seconds ahead into critical preview zones [2, 4]. Inspired by these human focus patterns, our research introduces the ’Focusing Sampling’ method, aimed at enhancing the detection and regression of distant lane boundary markings, a key challenge in high-speed autonomous driving. Additionally, ’Partial Field of View Evaluation’ is proposed to improve accuracy assessments in real-world scenarios, emphasizing the forward road sections align with driver focus.
@@
In the realm of lane detection, prevalent models such as CLRNet [5] and GANet [6] rely on uniform multi-scale feature pyramid network (FPN) architectures. These models, however, do not adequately capture the global spatial context, essential for detailing key road aspects for safe navigation. Even though transformers are known for their global semantic processing capabilities [7], their efficiency diminishes when identifying slender, elongated lane markings, where cardinal cues are sparse. Lane boundaries also display substantial scale variations and are affected by diverse lighting and surface conditions [8]. To overcome these challenges, our study introduces either positional or standard non-local blocks into the FPN, enriching it with extensive global context information. Furthermore, subsequent experiments demonstrate that ’Directional IoU Loss’ not only matches but can even exceed the benefits offered by positional non-local blocks.
This research contributes four innovations: (1) ’Focusing Sampling,’ a training strategy prioritizing small and distant lane details; (2) ’Partial Field of View Evaluation,’ new metrics for accuracy in forward road sections critical for real-world applications; (3) an enhanced FPN architecture that incorporates either positional non-local blocks or standard non-local blocks, depending on the requirement; (4) ’Directional IoU Loss,’ a novel regression loss that addresses directional discrepancies in distant lanes. FENetV1, employing positional non-local blocks, achieves state-of-the-art results on conventional metrics by concentrating on perspective-dependent semantics. On the other hand, FENetV2, which integrates coordinate modelling into the ’Directional IoU Loss’, excels in boundary localization accuracy for distant lanes. Although FENetV2 may underperform slightly in conventional metrics compared to FENetV1, its specialization in distant lane regression makes it more suited for practical navigation. In essence, this pioneering work advocates for the enhancement of regression-focused models and evaluations, employing targeted techniques that prioritize the representation and assessment of road details vital for accurate and safe autonomous lane detection.
@@ 2 Related works
Deep learning-based lane prediction methods can be categorized into semantic segmentation, anchor-based, and parameter-based approaches. Semantic segmentation methods like SCNN [9], SAD [10], and Curvelanes-NAS [11] use pixel-level predictions for accuracy but have high computational complexity. Anchor-based methods like CLRNet [5], GANet [6], E2E-LMD [12], IntRA-KD [13], UFLD/UFLDV2 [14, 15], CondLaneNet [16], and LaneATT [17] are fast but less accurate in complex scenarios. Parameter-based methods like ”Efficient Lane Detection via Curve Modeling” [18] use curve anchors for efficiency but lack optimal accuracy. We build on prior work to address limitations in accuracy and complexity.
@@ 3.1 Focusing Sampling
Motivation. Existing uniform point sampling strategies fail to properly weigh the visual perspective critical to robust lane detection. These methods treat all image regions equally, despite lanes containing more valuable semantic information at distant vanishing points, especially along curves [19]. To address the resulting deficiencies in understanding upcoming lanes, we propose a Focusing Sampling technique inspired by adept drivers’ visual attention directed far down the road. Like drivers who gaze ahead to foresee curves [1], Focusing Sampling emphasizes distant details while still examining nearby points. As demonstrated in Fig. 3, this novel approach properly accounts for complete lane geometry and handles complex turn and curve cases. Focusing Sampling is targeted to transform limitations of uniform sampling that risk losing data and semantics critical to lane prediction.
@@
where $a_{n}=\frac{1}{N_{sample}-1}*i$. $feature$ is the feature point number in the network. To produce focusing sample points in this range, a logarithmic transformation is applied to the feature distribution, mapping the original feature points to integer sample points. $H$ denotes the image height and $N_{sample}$ is the number of sampling points, while $a_{n}$ represents an arithmetic sequence generated within the range of 0 to 1 across $i$ data points. This logarithmic approach selects valuable semantic information. As an application of the logarithmic discretization introduces the possibility of repeated integer values, post-processing includes deduplication of the sample points. Complete details on the feature discretization process and resulting distribution are provided visually in the code and Appendix Fig. 8.
@@ 3.2 Positional Non-local Block and Position Enhanced FPN structure
Motivation. CLRNet principally employs the Feature Pyramid Network (FPN) architecture for lane detection [5]. While excelling at multi-scale feature extraction, FPN possesses certain limitations. Specifically, deeper network layers emphasize semantic representations, whereas shallow layers prioritize fine-grained details, culminating in suboptimal detection of diminutive targets. To augment FPN, we introduce a structure assimilating non-local blocks [20], inspired by enhancements in PANet and Mask R-CNN [21, 22]. This approach injects contextual clues into the network while preserving its depth and enhancing its multi-scale properties. Moreover, based on the distinctiveness of lane marking objectives, we propose a Position Enhanced FPN(PEFPN) module to actualize tighter integration between global semantics and lane coordinate modelling. As delineated in Fig. 2 Positional non-local blocks, and expanded upon Fig. 7 in Appendix A.1, position information is directly infused into multi-scale feature maps within the Enhanced FPN.
@@
This positional non-local block provides an enriched set of features encoding both global contexts as well as precise spatial locations. Furthermore, we apply the proposed Focusing Sampling technique on top of this architecture to selectively concentrate on lane segments during feature extraction. The samples provide strong cues to track lane coordinates and directionality. The FENetV1 incorporates PEFPN and Focusing Sampling that unifies refined coordinate modelling through direct coordinate injection with Focusing Sampling guided by lane positions. The ablation experiment in Appendix A.1 shows that the PEFPN represents a unique architecture designed according to the Focusing perspective for lane detection introduced in FENetV1.
@@ 3.3 Focusing Enhanced Network
Motivation. Within FENetV1, the formulation of PEFPN hinges upon the consolidated design of positional non-local blocks and Enhanced FPN architecture. Capitalizing on the distinctiveness of lane detection objectives, positional non-local blocks are devised based on coordinate map information, furnishing efficacious feature extraction. However, ablation experiments (elaborated in Appendix A.1) reveal that in the absence of Focusing Sampling, employing standard non-local blocks rather than positional non-local blocks conferred analogous performance. This validates that orientational clues from lane coordinates need not necessarily be injected at non-local locations. Therefore, guided by improving network efficiency, we conceived FENetV2 to further investigate Focusing Enhanced FPN(FEFPN) with standard non-local blocks and Directional IoU. Hence in devising FENetV2, we explore more efficient alternatives through standard non-local blocks, focusing enhancement modules, and Directional IoU formulations to preserve efficacy while enhancing computational thrift.
@@
Fig. 4: D-IoU Loss. D-IoU loss evaluates model lane prediction accuracy by comparing ground truth lanes to predicted lanes. It integrates the IoU between extended lane segments from sampled points $x_{i}$ along the lane. This provides a directional, localized loss metric accounting for prediction precision over the full lane geometry.
@@ 3.5 Training and Inference Details
Inference. To produce the final lane detection results, we apply non-maximum suppression (NMS) to remove overlapping predictions, leaving only optimal lanes [17]. We also evaluate an NMS-free approach by setting top-k to 1 during one-to-one assignments. These refinements extract the most accurate lanes.
@@ 3.6 Augmenting Evaluation via Partial Field of View
Existing evaluation metrics for lane detection models exhibit critical limitations aligning with real-world driving needs. Metrics like pixel-wise accuracy and standard mAP treat all pixels in lane marking predictions equally, despite significant differences in their actual safety impact. However, as noted previously, experienced drivers preferentially gaze at distant road regions, a strategy crucial for anticipating path geometry and necessary steering adjustments. To address this mismatch between current evaluations and necessary lane understanding for driving, we propose augmenting assessment through a Partial Field of View metric. This practical approach subdivides the lower image half into fractional views after preprocessing cropping of irrelevant upper content. As visualised in Fig. 5, by quantifying accuracy on key distal fraction views (the top 1/2 and top 1/3 fields), the metric better gauges precision for the most critical lane regions corresponding to human driver gaze. The Partial Field analysis provides a closer approximation to human perception. It constitutes an interpretable enhancement measuring lane detection aptitude through the practical lens of on-road gaze patterns.
@@ 4.1 Datasets
Table 1: FENet frameworks achieve state-of-the-art CULane performance with efficient GFlops. FENetV1 obtains the top mF1 and F1@75 scores, surpassing all methods including optimal CLRNet, showcasing advanced lane recognition. FENetV2 closely matches V1 metrics and we recommend it more for practical use, as analyzed in subsequent sections regarding its advantages.
@@ 4.4 Comparison with State-of-the-Art Results
Performance by Field of View. As Table 2 shows, compared to the state-of-the-art CLRNet model, FENetV2 demonstrates significantly improved curved and distant lane detection across the entire, the top half, and the top third field of views. Importantly, in the vast majority of driving scenarios, bend characteristics are mainly reflected from afar. Thus these gains primarily represent an enhanced perception of distant curved lanes, which is critical for safe maneuvering. Specifically, as shown by data marked with *, FENetV2 achieves 2.67, 5.66, and 6.01 higher mF1 scores for detecting challenging far-ahead curved lanes in that respective field of views. This highlights FENetV2’s particular strengths in precisely identifying tricky distant and curved lanes, enabled by its D-IoU loss function and FEFPN module. Some supporting example results are provided in Fig. 6 in Appendix A.1.
Although FENetV1 achieves slightly higher overall mF1 scores, we argue that FENetV2 is much more suitable and reliable for real-world self-driving systems due to its specialization for distant lane boundary regression. Accurately localizing lane lines far ahead is vital for timely vehicle control reactions, especially at faster speeds. In contrast, while FENetV1’s PEFPN exhibits strengths in general lane recognition, it has limitations in precision boundary localization. In summary, FENetV2 demonstrates the best real-world lane detection competency due to its specialization in distant lane regression, unlike prior state-of-the-art models. For practical autonomous navigation requiring split-second reactions, we strongly advocate for FENetV2 over existing methods and even the baseline FENetV1 model.

### v6

@@ 1 Introduction
This study highlights the difference between human visual focus during driving and the perspectives captured by 2D cameras. Fig. 1 shows that experienced drivers focus on distant road regions to anticipate path changes and steering needs, particularly around curves, looking ahead 1-2 seconds into crucial preview zones [1, 2, 3, 2, 4]. Inspired by these patterns, we introduce the ’Focusing Sampling’ method to better detect and regress distant lane boundaries, crucial for high-speed autonomous driving. We also propose the ’Partial Field of View Evaluation’ to enhance accuracy assessments in real-world scenarios by focusing on forward road sections that align with driver focus.
@@
In lane detection, widely-used models like CLRNet [5] and GANet [6] utilize standard multi-scale FPN architectures but fail to fully capture the necessary global spatial context for accurate road mapping crucial for safe navigation. Although transformers excel in global semantic processing, their effectiveness reduces in identifying thin, extended lane markings due to sparse cues [7]. Lane boundaries vary greatly in scale and are influenced by different lighting and surface conditions [8]. To address these issues, our study enhances the FPN by integrating either positional or standard non-local blocks, providing richer global context. Moreover, our experiments show that ’Directional IoU Loss’ equals or surpasses the advantages of positional non-local blocks.
This research introduces four key innovations: (1) ’Focusing Sampling,’ a training approach that emphasizes small and distant lane details; (2) ’Partial Field of View Evaluation,’ new metrics for accuracy in critical forward road sections; (3) an improved FPN architecture that includes either positional or standard non-local blocks; (4) ’Directional IoU Loss,’ a unique regression loss for correcting directional errors in distant lanes. FENetV1 uses positional non-local blocks for perspective-sensitive semantics, achieving top results with conventional metrics. Meanwhile, FENetV2, incorporating coordinate modeling with ’Directional IoU Loss,’ excels in precisely locating distant lane boundaries. Although it may slightly lag behind FENetV1 in traditional metrics, FENetV2’s focus on distant lane regression better suits real-world navigation. Overall, this work pushes for advances in regression-centric models and evaluations, focusing on techniques that enhance the depiction and assessment of critical road details for effective and safe autonomous lane detection.
@@ 2 Related works
Deep learning-based lane prediction methods fall into three categories: semantic segmentation, anchor-based, and parameter-based. Semantic segmentation approaches like SCNN, SAD, and Curvelanes-NAS offer pixel-level accuracy but are computationally demanding [9, 10, 11]. Anchor-based methods, including CLRNet, GANet, and others, are quicker but struggle with accuracy in complex scenarios [5, 6, 12, 13, 14, 15]. Parameter-based approaches, such as those using curve modeling, prioritize efficiency but compromise accuracy [16]. Our work aims to improve on these methods by addressing their accuracy and complexity limitations.
@@ 3.1 Focusing Sampling
Motivation. Uniform point sampling strategies in lane detection do not prioritize visual perspective, treating all image regions equally despite the importance of distant vanishing points, especially in curves [17]. To overcome these shortcomings, we introduce Focusing Sampling, inspired by how skilled drivers focus on distant road sections to anticipate curves [1]. This method emphasizes distant details while also considering nearby points, effectively capturing the full lane geometry and addressing complex turns and curves, as shown in Fig. 3. Focusing Sampling aims to improve upon uniform sampling by preserving critical data and semantics essential for accurate lane prediction.
@@
The term $a_{n}=\frac{1}{N_{sample}-1}*i$ represents an arithmetic sequence scaling from 0 to 1 across $i$ data points, where $N_{sample}$ is the number of sample points and $H$ is the image height. To generate focusing sample points, the feature point distribution undergoes a logarithmic transformation, converting feature points into integer sample values. This method emphasizes the extraction of key semantic information. Due to the possibility of repeated values from logarithmic discretization, deduplication is applied in post-processing. The feature discretization process and its output are detailed visually in the code and further illustrated in Appendix Fig. 8.
@@ 3.2 Positional Non-local Block and Position Enhanced FPN structure
Motivation. CLRNet uses the Feature Pyramid Network (FPN) architecture for lane detection [5], excelling in multi-scale feature extraction but facing challenges in detecting small targets due to its layer structure—deeper layers focus on semantics and shallower layers on fine details. To improve FPN, we incorporate non-local blocks [18], drawing on advancements from PANet and Mask R-CNN [19, 20], to enhance context awareness, depth, and multi-scale capabilities. Additionally, we introduce a Position Enhanced FPN (PEFPN) module tailored for lane detection, integrating global semantics with lane coordinate modeling. This integration is detailed in Appendix A.1 Fig. 2, showing direct infusion of position information into the FPN’s multi-scale feature maps.
@@
The positional non-local block in our architecture enhances feature extraction by encoding global context and precise spatial locations. We also integrate a Focusing Sampling technique that targets specific lane segments to better track lane coordinates and directions. FENetV1 combines this with PEFPN, which incorporates direct coordinate injection and Focusing Sampling aligned with lane positions, enhancing coordinate modeling. The ablation study in Appendix A.1 demonstrates that PEFPN is a uniquely designed architecture optimized for the focused detection of lanes in FENetV1.
@@ 3.3 Focusing Enhanced Network
Motivation. In FENetV1, the PEFPN framework combines positional non-local blocks with an Enhanced FPN architecture, utilizing coordinate maps for effective feature extraction in lane detection. Ablation studies (shown in Appendix A.1) indicate that without Focusing Sampling, standard non-local blocks perform similarly to positional ones, suggesting lane coordinates need not necessarily be injected at non-local locations. Thus, FENetV2 aims to boost network efficiency by incorporating standard non-local blocks and exploring Focusing Enhanced FPN (FEFPN) and Directional IoU. This version seeks to maintain effectiveness while improving computational efficiency.
@@ 3.4 Lane Directional Intersection over Union (D-IoU) Module
Fig. 4: D-IoU Loss. D-IoU loss assesses lane prediction accuracy by comparing predicted lanes to ground truth, using the IoU of extended lane segments from sampled points along the lane. This directional, localized loss metric accurately reflects predict precision across the entire lane geometry.
@@ 3.6 Augmenting Evaluation via Partial Field of View
Current metrics for evaluating lane detection models, like pixel-wise accuracy and standard mAP, fail to reflect the real-world driving needs, treating all lane marking pixels equally despite varying safety impacts. Notably, experienced drivers focus on distant road sections, vital for predicting path geometry and steering adjustments. To better align evaluations with actual driving requirements, we suggest using a Partial Field of View metric. This method divides the lower half of an image into segments after cropping out the top irrelevant area. As shown in Fig. 5, this metric evaluates accuracy in critical distant areas (top half and top third of the field), closely matching where drivers typically look, thus providing a more accurate reflection of lane detection performance through the perspective of human driving behavior.
@@ 4.1 Datasets
Table 1: FENet frameworks deliver top CULane performance with efficient GFlops usage. FENetV1 achieves the highest mF1 and F1@75 scores, outperforming all methods, including CLRNet, in lane recognition. FENetV2 performs comparably to V1, but is recommended for practical use, with its advantages detailed in subsequent sections.
@@ 4.4 Comparison with State-of-the-Art Results
Performance by Field of View. As shown in Table 2, FENetV2 outperforms the CLRNet model in detecting curved and distant lanes across all, the top half, and the top third fields of view. This improvement is crucial in driving scenarios where bends are primarily visible from afar, enhancing safe manoeuvring by better detecting distant curves. Specifically, FENetV2 achieves significantly higher mF1 scores of 2.67, 5.66, and 6.01 in these * data, underscoring its ability to accurately identify challenging distant and curved lanes. This performance is attributed to its D-IoU loss function and FEFPN module, with detailed examples provided in Fig. 6 in Appendix A.1.
Although FENetV1 scores slightly higher in overall mF1 metrics, FENetV2 is more effective and reliable for real-world autonomous driving due to its focus on distant lane boundary regression. This capability is crucial for accurate lane localization needed for quick vehicle control at high speeds. While FENetV1 excels in general lane detection, it lacks precision in boundary localization. In essence, FENetV2 stands out in practical lane detection performance by specializing in distant lane regression, making it preferable for autonomous navigation where immediate responses are critical. We recommend FENetV2 over other models, including FENetV1, for its superior real-world applicability.